# Square Functions Controlling Smoothness and Higher-Order Quantitative Rectifiability

### v1

@@ 5. Appendix
$\phi(x):=\theta(x)\left(ax_{i_{1}}^{n_{1}}...x_{i_{j}}^{n_{j}}\right)$
@@
$\left(\partial_{i_{1}}^{n_{1}}...\partial_{i_{j}}^{n_{j}}\right)\phi(x)\bigg{|}_{x=0}\neq 0$
@@
$\left(\partial_{i^{\prime}_{1}}^{n_{1}^{\prime}}...\partial_{i_{j^{\prime}}}^{n_{j}^{\prime}}\right)x_{i_{1}}^{n_{1}}...x_{i_{j}}^{n_{j}}\bigg{|}_{x=0}=0$
@@
$|\langle u,\varphi\rangle|\leq CQ(\varphi)$
@@
$|\langle b,\varphi\rangle|:=\left|\int b(x)\varphi(x)\,\mathrm{d}x\right|\leq CQ_{b}(\varphi)$
@@
$\int|b(x)\varphi(x)|\,\mathrm{d}x\lesssim\int\frac{|b(x)|}{1+|x|^{d+1}}\,\mathrm{d}x<\infty$
@@
$\int\frac{|b(x)-\langle b\rangle_{B_{1}(0)}|}{(1+|x|)^{n+1}}\leq c\|b\|_{*}$
@@
$\left|\int b(x)\varphi(x)\,\mathrm{d}x\right|=\left|\int(b(x)-\langle b\rangle_{B_{1}(0)})\varphi(x)\,\mathrm{d}x\right|\leq c(\varphi)\|b\|_{*}$
@@
$I_{\alpha}(b)(\varphi):=\int b(x)I_{\alpha}(\varphi)(x)\,\mathrm{d}x$
@@
$\lim_{\xi\rightarrow 0}|\xi|^{-\alpha}\hat{\varphi}(\xi)=0$
@@
$I_{\alpha}(\varphi)(x)=\int\varphi(y)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{n-\alpha}}\right)\,\mathrm{d}y$
So
$I_{\alpha}(b)(\varphi)=\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{n-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x$
@@
$\int\kern-3.39996pt\int\left|b_{k}(x)\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\right|\,\mathrm{d}y\,\mathrm{d}x\leq k\int|y|^{\alpha}|\varphi(y)|\,\mathrm{d}y<\infty$
@@
$\int b_{k}(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x=\int\kern-3.39996pt\int b_{k}(x)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y$
@@
$\left|\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x-\int\kern-3.39996pt\int b(x)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y\right|\\ \leq\left|\int(b(x)-b_{k}(x))\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x\right|+\left|\int\kern-3.39996pt\int(b(x)-b_{k}(x))\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y\right|\\ \leq\|b-b_{k}\|_{*}\|I_{\alpha}(\varphi)\|_{H^{1}}+\|b-b_{k}\|_{*}\int|y|^{\alpha}|\varphi(y)|\,\mathrm{d}y$
So, taking $k$ to infinity yields that
$\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x=\int\int b(x)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y$
@@
$I_{\alpha}(b)(x):=\int b(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|y|^{d-\alpha}}\right)\,\mathrm{d}y$
@@
$\partial_{i}f(\varphi)=-f(\partial_{i}\varphi),\,\,\,\varphi\in\mathcal{S}_{0}$
@@
$f(\varphi)=\langle b,I_{\alpha}(\varphi)\rangle$
@@
$\mathcal{R}_{i}b(\varphi):=\langle b,-\mathcal{R}_{i}\varphi\rangle,\,\,\,\varphi\in H_{0}^{1}$
@@
$\partial_{i}f(\varphi)=-\langle\tilde{b}_{i},I_{\alpha-1}(\varphi)\rangle=-\langle I_{\alpha-1}(\tilde{b}_{i}),\varphi\rangle$
@@
$\tilde{f}(x)=\int_{0}^{1}(\nabla f)(tx)\cdot x\,\mathrm{d}t$
@@
$=\sum_{i=1}^{d}\int_{0}^{1}\int_{\mathbb{R}^{d}}I_{\alpha-1}(-\tilde{b}_{i})(x)\frac{x_{i}}{t}\varphi\left(\frac{x}{t}\right)t^{-d}\,\mathrm{d}x\,\mathrm{d}t$
@@
$=\langle I_{\alpha}(b),\varphi\rangle$

### v2

@@ 5. Appendix
$\phi(x):=\theta(x)\left(ax_{i_{1}}^{n_{1}}...x_{i_{j}}^{n_{j}}\right).$
@@
$\left(\partial_{i_{1}}^{n_{1}}...\partial_{i_{j}}^{n_{j}}\right)\phi(x)\bigg{|}_{x=0}\neq 0.$
@@
$\left(\partial_{i^{\prime}_{1}}^{n_{1}^{\prime}}...\partial_{i_{j^{\prime}}}^{n_{j}^{\prime}}\right)x_{i_{1}}^{n_{1}}...x_{i_{j}}^{n_{j}}\bigg{|}_{x=0}=0.$
@@
$|\langle u,\varphi\rangle|\leq CQ(\varphi).$
@@
$|\langle b,\varphi\rangle|:=\left|\int b(x)\varphi(x)\,\mathrm{d}x\right|\leq CQ_{b}(\varphi).$
@@
$\int|b(x)\varphi(x)|\,\mathrm{d}x\lesssim\int\frac{|b(x)|}{1+|x|^{d+1}}\,\mathrm{d}x<\infty,$
@@
$\int\frac{|b(x)-\langle b\rangle_{B_{1}(0)}|}{(1+|x|)^{n+1}}\leq c\|b\|_{*}.$
@@
$\left|\int b(x)\varphi(x)\,\mathrm{d}x\right|=\left|\int(b(x)-\langle b\rangle_{B_{1}(0)})\varphi(x)\,\mathrm{d}x\right|\leq c(\varphi)\|b\|_{*}.$
@@
$I_{\alpha}(b)(\varphi):=\int b(x)I_{\alpha}(\varphi)(x)\,\mathrm{d}x.$
@@
$\lim_{\xi\rightarrow 0}|\xi|^{-\alpha}\hat{\varphi}(\xi)=0.$
@@
$I_{\alpha}(\varphi)(x)=\int\varphi(y)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{n-\alpha}}\right)\,\mathrm{d}y.$
So,
$I_{\alpha}(b)(\varphi)=\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{n-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x.$
@@
$\int\kern-3.39996pt\int\left|b_{k}(x)\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\right|\,\mathrm{d}y\,\mathrm{d}x\leq k\int|y|^{\alpha}|\varphi(y)|\,\mathrm{d}y<\infty.$
@@
$\int b_{k}(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x=\int\kern-3.39996pt\int b_{k}(x)\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y.$
@@
$\left|\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x-\int\kern-3.39996pt\int b(x)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y\right|\\ \leq\left|\int(b(x)-b_{k}(x))\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x\right|+\left|\int\kern-3.39996pt\int(b(x)-b_{k}(x))\left(\frac{1}{|x-y|^{n-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y\right|\\ \leq\|b-b_{k}\|_{*}\|I_{\alpha}(\varphi)\|_{H^{1}}+\|b-b_{k}\|_{*}\int|y|^{\alpha}|\varphi(y)|\,\mathrm{d}y.$
So, taking $k$ to infinity yields
$\int b(x)\int\varphi(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}y\,\mathrm{d}x=\int\int b(x)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|x|^{d-\alpha}}\right)\,\mathrm{d}x\varphi(y)\,\mathrm{d}y.$
@@
$I_{\alpha}(b)(x):=\int b(y)\left(\frac{1}{|x-y|^{d-\alpha}}-\frac{1}{|y|^{d-\alpha}}\right)\,\mathrm{d}y.$
@@
$\partial_{i}f(\varphi)=-f(\partial_{i}\varphi),\,\,\,\varphi\in\mathcal{S}_{0}.$
@@
$f(\varphi)=\langle b,I_{\alpha}(\varphi)\rangle.$
@@
$\mathcal{R}_{i}b(\varphi):=\langle b,-\mathcal{R}_{i}\varphi\rangle,\,\,\,\varphi\in H_{0}^{1}.$
@@
$\partial_{i}f(\varphi)=-\langle\tilde{b}_{i},I_{\alpha-1}(\varphi)\rangle=-\langle I_{\alpha-1}(\tilde{b}_{i}),\varphi\rangle.$
@@
$\tilde{f}(x)=\int_{0}^{1}(\nabla f)(tx)\cdot x\,\mathrm{d}t.$
@@
$=\sum_{i=1}^{d}\int_{0}^{1}\int_{\mathbb{R}^{d}}I_{\alpha-1}(-\tilde{b}_{i})(x)\frac{x_{i}}{t}\varphi\left(\frac{x}{t}\right)t^{-d}\,\mathrm{d}x\,\mathrm{d}t.$
@@
$=\langle I_{\alpha}(b),\varphi\rangle.$
@@
Acknowledgements: The author thanks Steve Hofmann for several helpful conversations. He also thanks the anonymous referee for several helpful suggestions which have imporved the manuscript.